# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ IV-A1 Hallucination from Data and Knowledge
The primary cause of hallucinations in LM agents stems from the biased or imperfect nature of the training data and knowledge employed to facilitate content generation. Specifically, i) annotation irrelevance in the collected dataset can lead to hallucinations in LM agents. For LLMs, source-reference divergence occurs due to heuristic data collection, implying that the selected sources may lack sufficient information for content generation [138]. For LVMs, a substantial amount of instruction data is synthesized using LLMs. However, generated instructions may not accurately correspond to the content depicted in the images due to the unreliability of generative models [139]. ii) The presence of distribution imbalance in the training data is another cause for hallucinations. The training dataset contains significantly more positive samples than negative ones, biasing LMs towards responding ‘yes’ in binary discrimination scenarios [139, 140]. iii) Duplicates within the training data may not be adequately filtered during pre-processing. Consequently, the trained LMs might tend to produce content that mimics the patterns in these duplicates [141]. iv) The training data and the knowledge, which can be inherently hallucinatory due to bias, outdated content, or fabrication [142]. LMs agents trained on these data are susceptible to replicate or even amplify hallucinations. Anomalous retrieval from the knowledge database can also introduce biases into the dependent knowledge during content generation process, thereby leading to the emergence of hallucinations in the generated content.
IV-A2 Hallucination from Training and Inference
@@
Parikh et al. [143] demonstrate that hallucinations can also occur even when the training dataset is nearly unbiased, and another key factor contributing to hallucinations is the presence of imperfections during LM training and inference processes. Specifically, i) defective encoder and decoder components in LM agents can result in hallucinations. The flawed representation learning in LM training increases the unexpected nature of the generation, thereby increasing the likelihood of hallucinated content [143]. Moreover, the sampling-based top-p decoding strategy has been demonstrated to be positively correlated with increased hallucination of generated content due to its inherent randomness [144]. ii) Parametric knowledge refers to LMs’ ability to utilize model parameters to memorize relevant information (e.g., knowledge) from training data, which can assist in improving the performance of downstream tasks [122]. Recent research has found that LMs prioritize parametric knowledge over user input during content generation [145], implying that biased parametric knowledge can result in a large amount of hallucinated information. iii) A problematic alignment process can also lead to hallucinations in LM agents. If the necessary prior knowledge is not adequately acquired during LM pre-training phase, the agent may produce hallucinated responses. Additionally, sycophancy, where the agent generates answers that align more with the user’s viewpoint than with accurate and credible information, can also contribute to hallucinations [146]. iv) The token generation strategy of LM agents may give rise to a phenomenon known as hallucination snowballing, wherein LM agent tends to persist in early errors for self-consistency rather than correcting them [147].
IV-A3 Hallucination from Multi-agent Interaction
@@
In multi-agent environments, interactions between LM agents can give rise to new hallucination threats. Specifically, i) conflicts between multiple LM agents may result in hallucinations in final responses. Due to the diverse objectives, strategies, and knowledge of individual agents, conflicting viewpoints or misinformation can inadvertently influence the final output when these agents engage in communication or collaboration. The potential adversarial manipulation may exacerbate the occurrence of conflicts, thereby leading to more severe hallucinations. ii) In multi-agent systems, the occurrence of a single agent’s hallucination can initiate a cascading effect, meaning that misinformation from one agent may be accepted and further propagated by others within the multi-agent network [41, 91]. Addressing hallucinations in multi-agent networks requires not only to correct inaccuracies at the individual agent level but also to effectively manage the exchange of information among agents, thereby preventing the spread of hallucinations.
IV-A4 Countermeasures to Hallucination
@@
Researchers have developed a suite of countermeasures targeting hallucination threats within LM agents. We outline these solutions alongside the recent advancements in research.
@@
  • Data sanitization represents an intuitive approach to mitigate hallucinations by excluding as much unreliable data as possible from training dataset. The manually-revised controlled dataset named ToTTo is validated to improve the factuality of results on table-to-text generation tasks [143]. Additionally, RefinedWeb, an automatically filtered and deduplicated web dataset, can mitigate hallucinations and improve LLM performance [142].
@@
  • Reinforcement learning prevents LMs from generating hallucinatory information by introducing external feedbacks and reward mechanisms to constrain models’ behavior, facilitating LM agents to refrain from QA beyond their capability instead of fabricating untruthful responses [148, 149]. For example, GPT-4 collects the unfactual data annotated by users and generates synthetic closed-domain hallucinated synthetic data to train the reward model, reducing the model’s tendency to hallucinate [6].
@@
  • Hallucination detection in generated content and subsequently producing samples devoid of such inaccuracies through re-generation is a promising strategy aimed at minimizing hallucination. Techniques such as SelfcheckGPT [150] and INSIDE [151] involve assessing the consistency across multiple generated responses to identify the occurrence of hallucinations.
@@
  • Truthful external knowledge & RAG techniques serve as convenient and efficient solutions to facilitate truthful content generation [15, 149]. They concatenate the external knowledge and user input to establish the context knowledge, and then use them to query LM agents to effectively reduce hallucinations in responses [152, 153].
@@
  • Instruction Tuning by optimizing instructions or fine-tuning LMs on a robust instruction dataset is an effective approach to mitigate hallucinations. SYNTRA [154] evaluates hallucinations on synthesis tasks and optimizes transferred instructions to reduce hallucinations on downstream tasks. Moreover, Liu et al. [139] mitigate hallucinations in LVMs by constructing a robust instruction dataset LRV-Instuction and fine-tuning LVMs on it.
@@
  • Post-processing methods can correct hallucinations in LM responses [155, 156]. RARR [155] employs search engines to collect related external knowledge and rectifies the generated content from multiple perspectives. LURE [156] reconstructs descriptions with less hallucinations by correcting object hallucinations during the post-process.
@@
  • Model architecture optimization is also an effective method to mitigate hallucinations. Multi-branch decoder [157] and uncertainty-aware decoder [158] are two examples of modified decoder structures to mitigate hallucinations.
@@
IV-B Adversarial Attack
@@
Figure 17: Illustration of adversarial attack to LM agents.
@@
For traditional AI models, an adversarial attack involves an adversary subtly manipulating the input by adding imperceptible perturbations, causing the AI model to produce outputs that deviate from the expected results. In the realm of LM agents, there are two types of adversarial attacks: adversarial input attacks and prompt hacking attacks, as depicted in Fig. 17. This subsection provides an in-depth overview of state-of-the-arts on these two types of attacks and the corresponding countermeasures.
IV-B1 Adversarial Input Attack
@@
Adversarial input attacks are analogous to traditional generative adversarial attacks, where the attacker degrades the accuracy of generated content in LM agents by manipulating input instructions, as depicted in Fig. 17(a). For example, an adversarial input attacker may subtly alter the text of an input news article to mislead the content summary agent, resulting in an incorrect or nonsensical summary. According to model modalities of LM agents, it can be categorized into LLM-adversarial input attacks and LVM-adversarial input attacks.
@@
  • LLM-adversarial input attack refers to the adversarial attack on LLMs that perturbs the input text into the semantic equivalent adversarial text, aiming to confuse LLMs to generate erroneous contents or adversary-desired contents. Zhuo et al. [31] demonstrate that CODEX, a pre-trained LLM capable of executing prompt-based semantic parsing on codes, is vulnerable to adversarial inputs, especially those generated by sentence-level perturbations. Xu et al. [159] propose PromptAttack, which converts adversarial textual attacks into an attack prompt, thereby causing the target LLM to generate adversarial examples to fool itself. Shi et al. [160] show that LLMs can be easily distracted by adding a small amount of irrelevant information to inputs, which significantly reduces the accuracy of their responses. Liu et al. [161] highlight that current LLMs do not involve social interactions during their training process. Consequently, the oversight leads to poor generalization ability in unfamiliar cases and a lack of robustness against adversarial attacks.
@@
  • LVM-adversarial input attack refers to the adversarial attack on LVMs (e.g., CLIP [69] and MiniGPT-4 [162]) that adds carefully crafted imperceptible adversarial perturbations to vulnerable modalities in the joint input prompts (i.e., visual, textual, or both), thereby compromising corresponding outputs. Du et al. [163] introduce a new adversarial attack named auto-attack on text-to-image models by adding small perturbations generated by gradient optimization to text prompts, resulting in the fusion of main subjects with unrelated categories or even their complete disappearance in generated images. The transferability of adversarial input attacks in LVMs has been extensively studied [164, 165]. Wang et al. [164] introduce the Transferable Multi-Modal (TMM) attack, which integrates attention-directed feature perturbation and orthogonal-guided feature heterogenization to generate transferable adversarial examples against LVMs; and Luo et al. [165] develop the Cross-Prompt Attack (CroPA), which leverages the learnable prompts to generate transferable adversarial images across LVMs. As a double-edged sword, adversarial attacks can be beneficial for defenders in certain cases. Liang et al. [166] leverage these attacks in the context of diffusion models to inhibit the learning, imitation, and replication of images by such models, thereby protecting intellectual property rights of artists.
IV-B2 Prompt Hacking Attack
@@
As depicted in Fig. 17(b), prompt hacking attacks involve using specifically crafted input instructions to bypass the security constraints of LM agents, thereby generating harmful contents. For example, an adversary could manipulate instructions given to a programming assistant agent to produce malicious code for ransomware. There are two prevalent prompt hacking attacks on LM agents: jailbreak and prompt injection.
@@
  • Jailbreak: LM agents typically achieve inherent predefined rule restrictions through model alignment techniques, thereby preventing the generation of harmful or malicious content. Jailbreak refers to adversaries designing particular prompts that exploit model vulnerabilities, bypassing the content generation rules, and enabling LM agents to generate harmful or malicious content [167, 168, 169, 170]. Yu et al. [167] evaluate hundreds of jailbreak prompts on GPT-3.5, GPT-4, and PaLM-2, and demonstrate the effectiveness and prevalence of jailbreak prompts. Yang et al. [168] propose an automated jailbreak attack framework named SneakyPrompt, which successfully jailbreaks DALL-E 2. SneakyPrompt effectively bypasses the safety filters and enable generation of Not-Safe-For-Work (NSFW) images. Shen et al. [169] perform a comprehensive measurement study on in-the-wild jailbreak prompts, identifying two long-term jailbreak prompts can achieve 99% attack success rates on GPT-3.5 and GPT-4. Deng et al. [170] introduce an end-to-end jailbreak framework named MASTERKEY, which utilizes the reverse engineering to uncover the internal jailbreak defense mechanisms of LLMs and leverages a fine-tuned LLM to automatically generate jailbreak prompts.
@@
IV-B3 Countermeasures to Adversarial Attacks
@@
  • Adversarial training aims to enhance AI model’s robustness in the input space by incorporating adversarial examples into the training data. Bespalov et al. [174] demonstrate that basic adversarial training can significantly improve the robustness of toxicity language predictors. Cheng et al. introduce AdvAug [175], a novel adversarial augmentation method for Neural Machine Translation (NMT) to enhance translation performance.
  • Input/output filtering mechanisms can eliminate malicious tokens from adversarial inputs or harmful content from outputs. Kumar et al. propose the erase-and-check method [176], which leverages another LLM as a safety filter to remove malicious tokens in the user input. Phute et al. [177] present an LLM self-examination defense approach, where an extra LLM is utilized to evaluate whether the responses are generated by adversarial prompts. Zeng et al. [178] propose AutoDefense, a multi-agent framework to defend against jailbreak attacks by filtering harmful LLM responses without impacting user inputs. AutoDefense divides the defense task into sub-tasks, leveraging LLM agents based on AutoGen [89] to handle each part independently. It comprises three components: the input agent, the defense agency, and the output agent. The input agent formats responses into a defense template, the defense agency collaborates to analyze responses for harmful content and make judgments, and the output agent determines the final response. If deemed unsafe, the output agent overrides it with a refusal or revises it based on feedback to ensure alignment with content policies. Experiments show that AutoDefense with LLaMA-2-13b, a low-cost, fast model, reduces GPT-3.5’s attack success rate (ASR) from 55.74% to 7.95%, achieving 92.91% defense accuracy.
@@
IV-C Poisoning & Backdoor Attack

### v2

@@ IV-C1 Hallucination from Data and Knowledge
The primary cause of hallucinations in LM agents stems from the biased or imperfect nature of the training data and knowledge employed for content generation. Specifically, i) annotation irrelevance in the collected dataset can lead to hallucinations in LM agents. For LLMs, source-reference divergence occurs due to heuristic data collection, implying that the selected sources may lack sufficient information for content generation [122]. For LVMs, a substantial amount of instruction data is synthesized using LLMs. However, generated instructions may not accurately correspond to the content depicted in the images due to the unreliability of generative models [123]. ii) The presence of distribution imbalance in the training data is another cause for hallucinations. The training dataset contains significantly more positive samples than negative ones, biasing LMs towards responding “yes” in binary discrimination scenarios [123, 124]. iii) Duplicates within the training data may not be adequately filtered during pre-processing. Consequently, the trained LMs might tend to produce content that mimics the patterns in these duplicates [125]. iv) The training data and the knowledge, which can be inherently hallucinatory due to bias, outdated content, or fabrication [126]. LMs agents trained on these data are susceptible to replicate or even amplify hallucinations. Anomalous retrieval from the knowledge base can also introduce biases into the dependent knowledge during content generation process, thereby leading to the emergence of hallucinations in the generated content.
IV-C2 Hallucination from Training and Inference
@@
Parikh et al. [127] demonstrate that hallucinations can occur even when the training dataset is nearly unbiased, and another key factor contributing to hallucinations is the presence of imperfections during LM training and inference processes. Specifically, i) defective encoder and decoder components in LM agents can result in hallucinations. The flawed representation learning in LM training increases the unexpected nature of the generation, thereby increasing the likelihood of hallucinated content [127]. Moreover, the sampling-based top-p decoding strategy has been demonstrated to be positively correlated with increased hallucination of generated content due to its inherent randomness [128]. ii) Parametric knowledge refers to LMs’ ability to utilize model parameters to memorize relevant information (e.g., knowledge) from training data, which can assist in improving the performance of downstream tasks [112]. Recent research has found that LMs prioritize parametric knowledge over user input during content generation [129], implying that biased parametric knowledge can result in a large amount of hallucinated information. iii) A problematic alignment process can also lead to hallucinations in LM agents. If the necessary prior knowledge is not adequately acquired during LM pre-training phase, the agent may produce hallucinated responses. Additionally, sycophancy, where the agent generates answers that align more with the user’s viewpoint than with accurate and credible information, can also contribute to hallucinations [130]. iv) The token generation strategy of LM agents may give rise to a phenomenon known as hallucination snowballing, wherein LM agent tends to persist in early errors for self-consistency rather than correcting them [131].
IV-C3 Hallucination from Multi-agent Interaction
@@
In multi-agent environments, interactions between LM agents can give rise to new hallucination threats. Specifically, i) conflicts between multiple LM agents may result in hallucinations in final responses. Due to the diverse objectives, strategies, and knowledge of individual agents, conflicting viewpoints or misinformation can inadvertently influence the final output when these agents engage in communication or collaboration. The potential adversarial manipulation may exacerbate the occurrence of conflicts, thereby leading to more severe hallucinations. ii) In multi-agent systems, the occurrence of a single agent’s hallucination can initiate a cascading effect, meaning that misinformation from one agent may be accepted and further propagated by others within the LM agent network [44, 68], especially under vertical collaboration paradigms. Addressing hallucinations in LM agent networks requires not only correcting inaccuracies at the individual agent level but also effectively managing the exchange of information among agents to prevent the spread of hallucinations.
IV-C4 Countermeasures to Hallucination in LM Agent Networks
@@
Researchers have developed a suite of countermeasures targeting hallucination threats within LM agents.
@@
  • Data sanitization represents an intuitive approach to mitigate hallucinations by excluding as much unreliable data as possible from training dataset. The manually-revised controlled dataset named ToTTo is validated to improve the factuality of results on table-to-text generation tasks [127]. Additionally, RefinedWeb is an automatically filtered and deduplicated web dataset, which can mitigate hallucinations and improve LLM performance [126].
@@
  • Reinforcement learning prevents LMs from generating hallucinatory information by introducing external feedbacks and reward mechanisms to constrain LM agents’ behavior, facilitating LM agents to refrain from QA beyond their capability instead of fabricating untruthful responses [132, 77]. For instance, GPT-4 collects the unfactual data annotated by users and generates synthetic closed-domain hallucinated synthetic data to train the reward model, reducing the model’s tendency to hallucinate [7].
@@
  • Hallucination detection is a promising strategy which detects hallucinations in generated content and subsequently regenerates samples to eliminate inaccuracies. Techniques such as SelfcheckGPT [133] and INSIDE [134] involve assessing the consistency across multiple generated responses to identify the occurrence of hallucinations.
@@
  • Truthful external knowledge & RAG techniques can facilitate truthful content generation [58, 77]. They concatenate external knowledge with user input to establish contextual knowledge, which is then used to query LM agents to reduce hallucinations in generated responses [135].
@@
  • Instruction tuning is an effective approach to mitigate hallucinations, by optimizing instructions or fine-tuning LMs on a robust instruction dataset. SYNTRA [136] evaluates hallucinations on synthesis tasks and optimizes transferred instructions to reduce hallucinations on downstream tasks. Moreover, Liu et al. [123] mitigate hallucinations in LVMs by constructing a robust instruction dataset named LRV-Instuction and fine-tuning LVMs on it.
@@
  • Post-processing techniques can help correct hallucinations in LM responses [137, 138]. For instance, RARR [137] utilizes search engines to gather relevant external knowledge and rectifies generated content from multiple perspectives. Meanwhile, LURE [138] reduces hallucinations by correcting object-level inaccuracies during the post-processing phase.
@@
  • Model architecture optimization is effective to mitigate hallucinations. Multi-branch decoder [139] and uncertainty-aware decoder [140] are two examples of modified decoder structures to mitigate hallucinations.
@@
IV-D Adversarial Attacks in LM Agent Networks
@@
Figure 24: Illustration of adversarial attacks to LM agents.
@@
For traditional AI models, an adversarial attack involves an adversary subtly manipulating the input by adding imperceptible perturbations, causing the AI model to produce outputs that deviate from the expected results. In the realm of LM agents, there are two types of adversarial attacks: adversarial input attacks and prompt hacking attacks, as depicted in Fig. 24.
IV-D1 Adversarial Input Attack
@@
Adversarial input attacks are analogous to traditional generative adversarial attacks, where the attacker degrades the accuracy of generated content in LM agents by manipulating input instructions, as depicted in Fig. 24(a). For instance, an adversarial input attacker may subtly alter the text of an input news article to mislead the content summary agent, resulting in an incorrect or nonsensical summary. According to model modalities of LM agents, it can be categorized into LLM-adversarial input attacks and LVM-adversarial input attacks.
@@
  • LLM-adversarial input attacks refers to the adversarial attack on LLMs that perturbs the input text into the semantic equivalent adversarial text, aiming to confuse LLMs to generate erroneous contents or adversary-desired contents. Zhuo et al. [32] demonstrate that CODEX, a pretrained LLM capable of executing prompt-based semantic parsing on codes, is vulnerable to adversarial inputs, especially those generated by sentence-level perturbations. Xu et al. [141] propose PromptAttack, which converts adversarial textual attacks into an attack prompt, thereby causing the target LLM to generate adversarial examples to fool itself. Shi et al. [142] show that LLMs can be easily distracted by adding a small amount of irrelevant information to inputs, which significantly reduces the accuracy of their responses. Liu et al. [143] highlight that current LLMs do not involve social interactions during their training process. Consequently, the oversight leads to poor generalization ability in unfamiliar cases and a lack of robustness against adversarial attacks.
@@
  • LVM-adversarial input attack refers to the adversarial attack on LVMs (e.g., CLIP and MiniGPT-4) that adds carefully crafted imperceptible adversarial perturbations to vulnerable modalities in the joint input prompts (i.e., visual, textual, or both), thereby compromising corresponding outputs. Du et al. [144] introduce a new adversarial attack named auto-attack on text-to-image models by adding small perturbations generated by gradient optimization to text prompts, resulting in the fusion of main subjects with unrelated categories or even their complete disappearance in generated images. The transferability of adversarial input attacks in LVMs has been extensively studied [145, 146]. Wang et al. [145] introduce the transferable multi-modal (TMM) attack, which integrates attention-directed feature perturbation and orthogonal-guided feature heterogenization to generate transferable adversarial examples against LVMs. Luo et al. [146] develop the cross-prompt attack (CroPA), which leverages the learnable prompts to generate transferable adversarial images across LVMs. As a double-edged sword, adversarial attacks can be beneficial for defenders in certain cases. For instance, Liang et al. [147] leverage these attacks in the context of diffusion models to inhibit the learning, imitation, and replication of images by such models, thereby protecting intellectual property rights of artists.
IV-D2 Prompt Hacking Attacks
@@
As depicted in Fig. 24(b), prompt hacking attacks involve using specifically crafted input instructions to bypass the security constraints of LM agents, thereby generating harmful contents. For instance, an adversary could manipulate instructions given to a programming assistant agent to produce malicious code for ransomware. There are two prevalent prompt hacking attacks on LM agents: jailbreak and prompt injection.
@@
  • Jailbreak: LM agents typically achieve inherent predefined rule restrictions through model alignment techniques, thereby preventing the generation of harmful or malicious content. Jailbreak refers to adversaries designing particular prompts that exploit model vulnerabilities, bypassing the content generation rules, and enabling LM agents to generate harmful or malicious content [148, 149, 150, 151]. Yu et al. [148] evaluate hundreds of jailbreak prompts on GPT-3.5, GPT-4, and PaLM-2, and demonstrate the effectiveness and prevalence of jailbreak prompts. Yang et al. [149] propose an automated jailbreak attack framework named SneakyPrompt, which successfully jailbreaks DALL-E 2. SneakyPrompt effectively bypasses the safety filters and enable generation of not-safe-for-work (NSFW) images. Shen et al. [150] perform a comprehensive measurement study on in-the-wild jailbreak prompts, identifying two long-term jailbreak prompts can achieve 99% attack success rates on GPT-3.5 and GPT-4. Deng et al. [151] introduce an end-to-end jailbreak framework named MASTERKEY, which utilizes the reverse engineering to uncover the internal jailbreak defense mechanisms of LLMs and leverages a fine-tuned LLM to automatically generate jailbreak prompts. In practical environments, jailbreak poses a more significant threat to embodied LM agents. Zhang et al. [152] identify three risks associated with embodied AI agents based on LLMs: cascading vulnerability propagation, cross-domain safety misalignment, and conceptual deception challenges. The first risk involves exploiting inherent jailbreak vulnerabilities in LLMs to induce harmful behaviors. The second arises from the mismatch between the language output and action output spaces of embodied LLM agents, allowing for the rejection of harmful behavior linguistically while still executing harmful actions. The underlying reason of this risk may stem from the inherent differences in aligned data. Currently, there is a wealth of aligned data within the language space, including a significant amount of ethically sound texts. In contrast, aligned data oriented towards action outputs is relatively scarce. The third risk stems from the current lack of reasoning capabilities in embodied LLM agents, making them susceptible to conceptual deception by attackers, which can result in seemingly harmless actions leading to malicious outcomes. For instance, an embodied LLM agent may decline a direct command to “poison the person”, but follow a series of inducible instructions that lead to the same result, such as “put the poison in the person’s mouth”.
@@
IV-D3 Countermeasures to Adversarial Attacks in LM Agent Networks
@@
  • Adversarial training aims to enhance LM’s robustness in the input space by incorporating adversarial examples into the training data. Bespalov et al. [156] demonstrate that basic adversarial training can significantly improve the robustness of toxicity language predictors. Cheng et al. introduce AdvAug [157], a novel adversarial augmentation method for neural machine translation (NMT) to enhance translation performance.
  • Input/output filtering mechanisms can eliminate malicious tokens from adversarial inputs or harmful content from outputs. Kumar et al. [158] propose the erase-and-check method, which leverages another LLM as a safety filter to remove malicious tokens in the user input. Phute et al. [159] present an LLM self-examination defense approach, where an extra LLM is utilized to evaluate whether the responses are generated by adversarial prompts. Zeng et al. [160] propose AutoDefense, a multi-agent framework to defend against jailbreak attacks by filtering harmful LLM responses without impacting user inputs. AutoDefense divides the defense task into sub-tasks, leveraging LLM agents based on AutoGen [13] to handle each part independently. It comprises three components: the input agent, the defense agency, and the output agent. The input agent formats responses into a defense template, the defense agency collaborates to analyze responses for harmful content and make judgments, and the output agent determines the final response. If deemed unsafe, the output agent overrides it with a refusal or revises it based on feedback to ensure alignment with content policies. Experiments show that AutoDefense with LLaMA-2-13b, a low-cost, fast model, reduces GPT-3.5’s attack success rate (ASR) from 55.74% to 7.95%, achieving 92.91% defense accuracy.
@@
IV-E Poisoning & Backdoor Attacks in LM Agent Networks